OpenStudy (anonymous):

Hi, I have 5 questions that I need help with. 1. For the following system of equations, what is the x-value of the solution? -x+2y=6 6y=x+18 (Choices) -18 0 -6 3

OpenStudy (dayakar):

-x + 2y = 6------(1) -x + 6y =18 -----(2) multiply (1) * 3 and subtract (1) from (2)

OpenStudy (anonymous):

I'm still lost...

OpenStudy (dayakar):

multiply (1) both side with 3 what do u get

OpenStudy (anonymous):

This is my worst subject...

OpenStudy (dayakar):

3(-x +2y) = 3 * 6 can u do it now

OpenStudy (anonymous):

No

OpenStudy (dayakar):

i will give u example 2(x+ 3) = 2* 4 2*x + 2*3 = 8 do it like this